A string is stretched and fastened to two points distance l apart. Motion is started by displacing the string into the form y=k(lx-x^2),from which it is released at time t=0.write the governing differential equation and initial and boundary condition for this problem
1
Expert's answer
2020-10-06T18:50:35-0400

The displacement y(x,t) is given by the equation,


"\frac{d^2y}{dt^2}=a^2\frac{d^2y}{dt^2}"


The boundary condition are:-

i ) y(0,t)= 0 , for t"\ge0"

ii ) y(l,t)= 0 , for t"\ge0"


iii) y(x,0)= 0 , for "0\le x\le l"


iv) "[\frac{dy}{dt}]_{t=0}=kx(l-x), for 0\le x\le l"